# Gradient Descent for Noisy Optimization

### v3

@@ 1.1 Set-up and problem formulation
Finally, we note that Fehrman et al., (2020) study SG algorithms under the assumption that $F$ is locally $L$-smooth and allowing $\mathbb{E}[\|\nabla_{\theta}f(\theta,Z)\|^{2}]$ to be unbounded but, informally speaking, this reference only shows that SG will converge to $\theta_{\star}$ if the starting value of the algorithm is sufficiently close to that point.
Consequently, as illustrated by Propositions 1-2, even for seemingly simple noisy optimization problems on $\mathbb{R}^{d}$, where $F$ is strictly convex and $\mathbb{E}[\|\nabla_{\theta}f(\theta,Z)\|^{2}]$ is finite for all $\theta\in\mathbb{R}^{d}$, SG is not guaranteed to converge.
@@ 1.2 Contribution of the paper
More precisely, we assume throughout this work that there is a computational cost $C_{\mathrm{grad}}\in\mathbb{N}$ for computing $\nabla_{\theta}f(\theta,z)$ and a computational cost $C_{\mathrm{eval}}\in\mathbb{N}$ for computing $f(\theta,z)$ for a pair $(\theta,z)\in\mathbb{R}^{d}\times\mathsf{Z}$. Then, we first show that, given the available computational budget $B\in\mathbb{N}$, by applying GD-BLS to minimize the function $F_{n_{1}(B)}:=\frac{1}{n_{1}(B)}\sum_{i=1}^{n_{1}(B)}f(\cdot,Z_{i})$ we obtain a consistent estimator $\hat{\theta}_{1,B}$ of $\theta_{\star}$. The convergence rate of this approach, which is particular type of sample average approximation algorithms (see e.g. Kim et al.,, 2015), depends on the choice of the function $n_{1}$ and, assuming for now that $\mathbb{E}[\|\nabla_{\theta}f(\theta_{\star},Z)\|^{2}]<\infty$, the upper bound we obtain for this rate is minimized when $n_{1}(B)\approx B^{1/2}$, in which case $\|\hat{\theta}_{1,B}-\theta_{\star}\|=\mathcal{O}_{\mathbb{P}}(B^{-0.25})$. Remark that the $B^{-0.25}$ convergence rate obtained for $\hat{\theta}_{1,B}$ is slower than the standard $B^{-0.5}$ convergence rate.
@@
In some applications the variance of the noisy gradient is not finite, and in this work the procedure described above is studied under the weaker assumption that $\mathbb{E}[\|\nabla_{\theta}f(\theta_{\star},Z)\|^{1+\alpha}]<\infty$ for some $\alpha\in(0,1]$. In this scenario, the user needs to choose a noise parameter $\alpha^{\prime}\in(0,1]$ and a $\delta\in(2\alpha^{\prime}/(1+3\alpha^{\prime}),1)$, and the proposed approach for computing $\theta_{\star}$ is proved to be consistent for any $\alpha^{\prime}\in(0,1]$. In addition, if $\alpha^{\prime}=\alpha$ then its convergence rate for computing $\theta_{\star}$ is shown to be $B^{-\frac{\alpha}{1+\alpha}(1-\delta^{J})}$. We recall the reader that $B^{-\frac{\alpha}{1+\alpha}}$ is the best possible rate we can have for a first-order stochastic convex optimization algorithm under the assumption that $\mathbb{E}[\|\nabla_{\theta}f(\theta_{\star},Z)\|^{1+\alpha}]<\infty$ (see Vural et al.,, 2022, and references therein). To contrast these results with the known properties of SG we note that, recently, Nguyen et al., (2024) showed that clipped SG achieves the optimal $B^{-\frac{\alpha}{1+\alpha}}$ rate, assuming that $\alpha$ is known. However, the results obtained in this reference rely on the strong assumption that $\sup_{\theta\in\mathbb{R}^{d}}\mathbb{E}[\|\nabla_{\theta}f(\theta,Z)\|^{1+\alpha}]<\infty$ and require the function $F$ to be $L$-smooth. The behaviour of standard SG algorithms under infinite noise variance is studied in Wang et al., (2021), but a very particular structure on the noisy gradient $\nabla_{\theta}f(\theta,Z)$ is imposed in addition to an unusual smoothness condition on $F$.
@@ 1.3 Connection with retrospective approximation
The proposed $J$-step procedure can be seen as a retrospective approximation (RA) algorithm with a fixed computational budget and using GD-BLS as deterministic optimizer (see. Kim et al.,, 2015, for a survey on RA methods). Indeed, given a sequence $(n^{\prime}_{j})_{j\geq 1}$ in $\mathbb{N}$ and a sequence $(\tau_{j}^{\prime})_{j\geq 1}$ in $(0,\infty)$, RA based on GD-BLS amounts to applying the above $J$-step procedure with $n_{j}(B)=n^{\prime}_{j}$ and $\tau_{j}(B)=\tau^{\prime}_{j}$ for all $B\in\mathbb{N}$ and $j\in\mathbb{N}$, and with $B=\infty$ (so that the optimization of $F_{n_{j}^{\prime}}$ is performed for all $j\in\{1,\dots,J\}$). Unlike in the proposed approach, the computational budget $B_{J}$ consumed by an RA algorithm is therefore random. The behaviour of $\mathbb{E}[B_{J}]$ and the convergence rate, as $J\rightarrow\infty$, of the resulting estimator $\theta_{J}$ of $\theta_{\star}$ have been recently studied by Newton et al., (2024) under general conditions on the optimization routine $\mathcal{R}$ used to optimise $F_{n^{\prime}_{j}}$ for all $j\in\{1,\dots,J\}$. However, applying Newton et al., (2024, Theorem 6) to the case where $\mathcal{R}$ is gradient descent, or GD-BLS, only shows that $\|\theta_{J}-\theta_{\star}\|$ converges to zero at the slow $\mathbb{E}[B_{J}]^{-1/4}$ rate. In addition, the results presented in this latter reference are obtained under conditions which seem hard to fulfil if $F$ is not (globally) $L$-smooth.
@@ 2.2 Assumptions
Consider the set-up of Proposition 1. Then, Assumptions A1-A2 are satisfied with $\alpha=2$.
@@ 2.3 Some implications of Assumptions A1-A2
The last lemma establishes that, under A1 and A2, by using the first $n$ elements of $(Z_{i})_{i\geq 1}$ we can learn $\theta_{\star}$ at rate $n^{-\frac{\alpha}{1+\alpha}}$, with $\alpha\in(0,1]$ as in A2. In addition to be used to establish our main results, this lemma will allow us to discuss the convergence rate of the proposed estimators.
@@ 3.2.2 Convergence result
$\|\hat{\theta}_{J,B}-\theta_{\star}\|=\mathcal{O}\Big{(}B^{-\frac{\alpha^{\prime}}{1+\alpha^{\prime}}(1-\delta^{J})}\Big{)}.$
@@ 4.1 Example of Proposition 1
As shown above in Proposition 3, for this optimization problem Assumptions A1-A2 hold with $\alpha=2$ and, to the best of our knowledge, SG is not guaranteed to converge when used to solve it (see Propositions 1-2). Simple computations show that the global minimizer of the resulting objective function $F$ is $\theta_{\star}=0$. For this example we let $\theta_{0}=1$ and $\alpha^{\prime}=2$, in which case Theorem 2 imposes to have $\delta\in(0.5,1)$.
@@ 4.3 Example of Proposition 1 with an extra heavy tailed noise
We let $\nu=1.501$ so that A2 holds for any $\alpha<1-\nu=0.501$. Remark that since $\mathbb{E}[W]=0$ it follows that the resulting objective function $F$ is the same as for the example of Proposition 1, and thus $\theta_{\star}=0$. It is also immediate to see that the conclusions of Proposition 1-2 remain valid for this example, and therefore to be best of our knowledge SG is not guaranteed to converge when applied to solve it. In addition, the results of Nguyen et al., (2024) for clipped SG do not apply to this optimization problem, notably because the function $F$ is not $L$-smooth (see Proposition 1). For this example we implement Algorithm 4 with $\theta_{0}=1$ and the results are presented for $\alpha^{\prime}=0.5$ and for $\alpha^{\prime}=1$. Remark that A2 holds with $\alpha=\alpha^{\prime}$ for the former value of $\alpha^{\prime}$ but not for the latter. Remark also that, by Theorem 2, we must have $\delta\in(0.4,1)$ when $\alpha^{\prime}=0.5$ and $\delta\in(0.5,1)$ when $\alpha^{\prime}=1$.
@@ A.2.6 Proof of Lemma 5
Remark that, by (19)-(20) and Lemma 4, and since $\theta_{\star}\in\mathring{\Theta}$, we have $\liminf_{B\rightarrow\infty}\mathbb{P}(\Omega_{\epsilon,n})\geq 1-\epsilon$ for all $\epsilon\in(0,1)$. Remark also that $\lambda_{\Theta}>0$ under A1 and by Lemma 2, and thus for all $n\geq 1$ and $\epsilon>0$ the function $F_{n}^{\omega}$ is strictly convex on $\Theta$ for all $\omega\in\Omega_{\epsilon,n}$.
Let $n\geq 1$, $\epsilon\in(0,1)$ and $\omega\in\Omega_{\epsilon,n}$, and note that $\nabla F^{\omega}_{n}(\hat{\theta}^{\omega}_{\Theta,{n}})=0$ since $F^{\omega}_{n}$ is strictly convex on $\Theta$ and $\hat{\theta}^{\omega}_{\Theta,{n}}\in\mathring{\Theta}$. Therefore, since $\theta_{\star}\in\mathrm{\Theta}$, it follows from Taylor’s theorem that there exists a $\check{\theta}^{\omega}_{n_{B}}\in\mathring{\Theta}$ such that
@@ A.2.8 Proof of Proposition 5
and remark that $x+v\nabla g(x)\in K$ for all $x\in\tilde{K}$ and $v\in(0,1]$. We now let $x\in\tilde{K}\subset K$ and $v\in(0,1]$. Then, since by (26) we have
@@ A.3.3 A second key property of Algorithm 3
Next, assume first that $T_{\theta_{0},B}^{\omega}>0$ and for all $t\in\{1,\dots,T_{\theta_{0},B}^{\omega}\}$ let $v_{t}^{\omega}$ be such that $\theta^{\omega}_{\theta_{0},B,t}=\theta^{\omega}_{\theta_{0},B,t-1}+v_{t}^{\omega}\nabla F_{n_{B}}(\theta^{\omega}_{\theta_{0},B,t-1})$. Then, by applying Proposition 6 with $g=F_{n_{B}}^{\omega}$, $K=\tilde{\Theta}_{j+1,x}$, $\tilde{K}=\Theta^{\prime}_{j+1,x}$ and $L_{K}=L_{j+1,x}$, it follows that $v_{t}^{\omega}\geq\beta/L_{j+1}$ for all $t\in\{1,\dots,T_{\theta_{0},B}^{\omega}\}$. From this result, we readily obtain that
$C^{\omega}_{\theta_{0},B,t}\leq D_{j}n_{B},\quad D_{j}=(\beta^{k_{j}}+1)(C_{\mathrm{eval}}+C_{\mathrm{grad}}),\quad k_{j}=\inf\{k\in\mathbb{N}:\,\beta^{k}<\beta/L_{j+1}\}.$

### v4

@@ 1.1 Set-up and problem formulation
Finally, we note that Fehrman et al., (2020) study SG algorithms under the assumption that $F$ is locally $L$-smooth and allowing $\mathbb{E}[\|\nabla_{\theta}f(\theta,Z)\|^{2}]$ to be unbounded but, informally speaking, this reference only shows that SG will converge to $\theta_{\star}$ if the starting value of the algorithm is sufficiently close to that point. Consequently, as illustrated by Propositions 1-2, even for seemingly simple noisy optimization problems on $\mathbb{R}^{d}$, where $F$ is strictly convex and $\mathbb{E}[\|\nabla_{\theta}f(\theta,Z)\|^{2}]$ is finite for all $\theta\in\mathbb{R}^{d}$, SG is not guaranteed to converge.
If above we focused our attention on SG methods it is worth mentioning that stochastic proximal point methods (see Asi and Duchi,, 2019, and references therein) is another popular set of tools for solving optimization problems as defined in (1). These algorithms have the advantage of having stronger theoretical properties than SG approaches but, on the other hand, they usually have the drawback to require, at each iteration, to use a numerical method to approximate the solution of an optimization problem. The Poisson regression example introduced in Proposition 1 provides a simple illustration of this claim: convergence results for the stochastic proximal point method applied to this problem have been obtained by Asi and Duchi, (2019) but, for this example, there is no explicit solution to the optimization problem that needs to be solved at each iteration of the algorithm. We finally note that the results derived in Asi and Duchi, (2019) in fact apply to a broader class of algorithms, called model-based stochastic proximal point methods. In these algorithms, the choice of the model determines the optimization problem that needs to be solved at each iteration, and Asi and Duchi, (2019, Example 3) use the Poisson regression example as a motivating example for what they call the truncated model. However, the authors do not show that the resulting algorithm is stable, that is that their key condition (C.iv) used to establish the stability of model-based stochastic proximal point methods is satisfied.
Consequently, as illustrated by the Poisson regression example of Proposition 1, they are seemingly simple noisy optimization problems on $\mathbb{R}^{d}$, where $F$ is strictly convex and $\mathbb{E}[\|\nabla_{\theta}f(\theta,Z)\|^{2}]$ is finite for all $\theta\in\mathbb{R}^{d}$, which cannot be solved (easily) with theoretical guarantees using existing methods.
@@ 1.2 Contribution of the paper
More precisely, we assume throughout this work that there is a computational cost $C_{\mathrm{grad}}\in\mathbb{N}$ for computing $\nabla_{\theta}f(\theta,z)$ and a computational cost $C_{\mathrm{eval}}\in\mathbb{N}$ for computing $f(\theta,z)$ for a pair $(\theta,z)\in\mathbb{R}^{d}\times\mathsf{Z}$. Then, we first show that, given the available computational budget $B\in\mathbb{N}$, by applying GD-BLS to minimize the function $F_{n_{1}(B)}:=\frac{1}{n_{1}(B)}\sum_{i=1}^{n_{1}(B)}f(\cdot,Z_{i})$ we obtain a consistent estimator $\hat{\theta}_{1,B}$ of $\theta_{\star}$. The convergence rate of this approach, which is particular type of sample average approximation algorithms (see e.g. Kim et al.,, 2015), depends on the choice of the function $n_{1}$ and, assuming for now that $\mathbb{E}[\|\nabla_{\theta}f(\theta_{\star},Z)\|^{2}]<\infty$, the upper bound we obtain for this rate is minimized when $n_{1}(B)\approx B^{1/2}$, in which case $\|\hat{\theta}_{1,B}-\theta_{\star}\|=\mathcal{O}_{\mathbb{P}}(B^{-0.25})$. Remark that the $B^{-0.25}$ convergence rate obtained for $\hat{\theta}_{1,B}$ is slower than the standard $B^{-0.5}$ convergence rate (see Proposition 4).
@@
In some applications the variance of the noisy gradient is not finite, and in this work the procedure described above is studied under the weaker assumption that $\mathbb{E}[\|\nabla_{\theta}f(\theta_{\star},Z)\|^{1+\alpha}]<\infty$ for some $\alpha\in(0,1]$. In this scenario, the user needs to choose a noise parameter $\alpha^{\prime}\in(0,1]$ and a $\delta\in(2\alpha^{\prime}/(1+3\alpha^{\prime}),1)$, and the proposed approach for computing $\theta_{\star}$ is proved to be consistent for any $\alpha^{\prime}\in(0,1]$. In addition, if $\alpha^{\prime}=\alpha$ then its convergence rate for computing $\theta_{\star}$ is shown to be $B^{-\frac{\alpha}{1+\alpha}(1-\delta^{J})}$. By contrast, in this context and as shown in Proposition 4, under the assumptions imposed on (1) in the next section $B^{-\frac{\alpha}{1+\alpha}}$ is the best possible convergence for estimating $\theta_{\star}$.
@@ 1.3 Connection with retrospective approximation
The proposed $J$-step procedure can be seen as a retrospective approximation (RA) algorithm with a fixed computational budget and using GD-BLS as deterministic optimizer (see. Kim et al.,, 2015, for a survey on RA methods). Indeed, given a sequence $(n^{\prime}_{j})_{j\geq 1}$ in $\mathbb{N}$ and a sequence $(\tau_{j}^{\prime})_{j\geq 1}$ in $(0,\infty)$, RA based on GD-BLS amounts to applying the above $J$-step procedure with $n_{j}(B)=n^{\prime}_{j}$ and $\tau_{j}(B)=\tau^{\prime}_{j}$ for all $B\in\mathbb{N}$ and $j\in\mathbb{N}$, and with $B=\infty$ (so that the optimization of $F_{n_{j}^{\prime}}$ is performed for all $j\in\{1,\dots,J\}$). Unlike in the proposed approach, the computational budget $B_{J}$ consumed by an RA algorithm is therefore random. The behaviour of $\mathbb{E}[B_{J}]$ and the convergence rate, as $J\rightarrow\infty$, of the resulting estimator $\theta_{J}$ of $\theta_{\star}$ have been recently studied by Newton et al., (2024) under general conditions on the optimization routine $\mathcal{R}$ used to optimise $F_{n^{\prime}_{j}}$ for all $j\in\{1,\dots,J\}$. However, the results presented in this reference are obtained under conditions which seem hard to fulfil if $F$ is not (globally) $L$-smooth.
@@ 2.2 Assumptions
Consider the set-up of Proposition 1. Then, Assumptions A1-A2 are satisfied with $\alpha=1$.
@@ 2.3 Some implications of Assumptions A1-A2
The following lemma establishes that, under A1 and A2, by using the first $n$ elements of $(Z_{i})_{i\geq 1}$ we can learn $\theta_{\star}$ at rate $n^{-\frac{\alpha}{1+\alpha}}$, with $\alpha\in(0,1]$ as in A2. In addition to be used to establish our main results, this lemma will allow us to discuss the convergence rate of the proposed estimators.
@@
Finally, the following proposition shows that, under Assumptions A1-A2 and in the context of this work (i.e. noisy optimization unded fixed computations budget) the best possible convergence rate for an estimator of $\theta_{\star}$ is $B^{-\frac{\alpha}{1+\alpha}}$, with $\alpha\in(0,1]$ as in A2.
Proposition 4.
Let $\alpha^{\prime}\in(0,1]$. Then, for any sequence $(\hat{\theta}_{B})_{B\geq 1}$ of estimators of $\theta_{\star}$ such that, for all $B\in\mathbb{N}$, $\hat{\theta}_{B}$ depends on $(Z_{i})_{i\geq 1}$ only through $k_{B}\in\mathbb{N}_{0}$ evaluations of $f(\cdot,\cdot)$ and $k_{B}^{\prime}\in\mathbb{N}_{0}$ evaluations of $\nabla f(\cdot,\cdot)$, where $k_{B}C_{\mathrm{eval}}+k_{B}^{\prime}C_{\mathrm{grad}}\leq B$, there is a noisy optimization problem (1) such that A1-A2 hold with $\alpha=\alpha^{\prime}$ and such that, for some constants $\delta\in(0,1)$ and $C\in(0,\infty)$, we have
$\limsup_{B\rightarrow\infty}\mathbb{P}\Big{(}\|\hat{\theta}_{B}-\theta_{\star}\|>CB^{-\frac{\alpha}{1+\alpha}}\Big{)}\geq\delta.$
@@ 3.2.2 Convergence result
$\|\hat{\theta}_{J,B}-\theta_{\star}\|=\mathcal{O}_{\mathbb{P}}\Big{(}B^{-\frac{\alpha^{\prime}}{1+\alpha^{\prime}}(1-\delta^{J})}\Big{)}.$
@@ 4.1 Example of Proposition 1
As shown above in Proposition 3, for this optimization problem Assumptions A1-A2 hold with $\alpha=1$ and, to the best of our knowledge, SG is not guaranteed to converge when used to solve it (see Propositions 1-2). Simple computations show that the global minimizer of the resulting objective function $F$ is $\theta_{\star}=0$. For this example we let $\theta_{0}=1$ and $\alpha^{\prime}=1$, in which case Theorem 2 imposes to have $\delta\in(0.5,1)$.
@@ 4.3 Example of Proposition 1 with an extra heavy tailed noise
We let $\nu=1.501$ so that A2 holds for any $\alpha<\nu-1=0.501$. Remark that since $\mathbb{E}[W]=0$ it follows that the resulting objective function $F$ is the same as for the example of Proposition 1, and thus $\theta_{\star}=0$. It is also immediate to see that the conclusions of Proposition 1-2 remain valid for this example, and therefore to be best of our knowledge SG is not guaranteed to converge when applied to solve it. In addition, the results of Nguyen et al., (2024) for clipped SG do not apply to this optimization problem, notably because the function $F$ is not $L$-smooth (see Proposition 1). For this example we implement Algorithm 4 with $\theta_{0}=1$ and the results are presented for $\alpha^{\prime}=0.5$ and for $\alpha^{\prime}=1$. Remark that A2 holds with $\alpha=\alpha^{\prime}$ for the former value of $\alpha^{\prime}$ but not for the latter. Remark also that, by Theorem 2, we must have $\delta\in(0.4,1)$ when $\alpha^{\prime}=0.5$ and $\delta\in(0.5,1)$ when $\alpha^{\prime}=1$.
@@ A.2.6 Proof of Lemma 5
Remark that, by (19)-(20) and Lemma 4, and since $\theta_{\star}\in\mathring{\Theta}$, we have $\liminf_{n\rightarrow\infty}\mathbb{P}(\Omega_{\epsilon,n})\geq 1-\epsilon$ for all $\epsilon\in(0,1)$. Remark also that $\lambda_{\Theta}>0$ under A1 and by Lemma 2, and thus for all $n\geq 1$ and $\epsilon>0$ the function $F_{n}^{\omega}$ is strictly convex on $\Theta$ for all $\omega\in\Omega_{\epsilon,n}$.
Let $n\geq 1$, $\epsilon\in(0,1)$ and $\omega\in\Omega_{\epsilon,n}$, and note that $\nabla F^{\omega}_{n}(\hat{\theta}^{\omega}_{\Theta,{n}})=0$ since $F^{\omega}_{n}$ is strictly convex on $\Theta$ and $\hat{\theta}^{\omega}_{\Theta,{n}}\in\mathring{\Theta}$. Therefore, since $\theta_{\star}\in\mathring{\Theta}$, it follows from Taylor’s theorem that there exists a $\check{\theta}^{\omega}_{n_{B}}\in\mathring{\Theta}$ such that
@@ A.2.7 Proof of Proposition 4
Proof of Proposition 4.
We consider the optimization problem (1) with $d=1$, with $\mathsf{Z}=\mathbb{R}$, with $f:\mathbb{R}\times\mathsf{Z}\rightarrow\mathbb{R}$ defined by $f(\theta,z)=\theta^{2}-2\theta z$, $(\theta,z)\in\mathbb{R}^{2}$ and with $Z\sim P$ for some probability distribution $P$. Then, it is direct to check that for any distribution $P$ such that $\mathbb{E}[|Z|^{1+\alpha^{\prime}}]<\infty$ Assumptions A1-A2 hold with $\alpha=\alpha^{\prime}$, and that we have $\theta_{\star}=\mathbb{E}[Z]$. For all $B\in\mathbb{N}$ let $n_{B}$ denote the number of distinct elements of $(Z_{i})_{i\geq 1}$ used to define $\hat{\theta}_{B}$. Then, by Devroye et al., (2016, Theorem 3.1), there exists a distribution $P$ of $Z$ such that $\mathbb{E}[|Z|^{1+\alpha^{\prime}}]<\infty$ and such that, for some constants $\delta\in(0,1)$ and $C\in(0,\infty)$, we have
$\limsup_{B\rightarrow\infty}\mathbb{P}\Big{(}|\hat{\theta}_{B}-\mathbb{E}[Z]|>Cn_{B}^{-\frac{\alpha}{1+\alpha}}\big{)}\geq\delta.$
The result of the proposition follows upon noting that $n_{B}\leq B$ for all $B\in\mathbb{N}$. ∎
A.2.8 Proof of Proposition 5
@@ A.2.9 Proof of Proposition 6
and remark that $x-v\nabla g(x)\in K$ for all $x\in\tilde{K}$ and $v\in(0,1]$. We now let $x\in\tilde{K}\subset K$ and $v\in(0,1]$. Then, since by (26) we have
@@ A.3.3 A second key property of Algorithm 3
Next, assume first that $T_{\theta_{0},B}^{\omega}>0$ and for all $t\in\{1,\dots,T_{\theta_{0},B}^{\omega}\}$ let $v_{t}^{\omega}$ be such that $\theta^{\omega}_{\theta_{0},B,t}=\theta^{\omega}_{\theta_{0},B,t-1}+v_{t}^{\omega}\nabla F_{n_{B}}(\theta^{\omega}_{\theta_{0},B,t-1})$. Then, by applying Proposition 7 with $g=F_{n_{B}}^{\omega}$, $K=\tilde{\Theta}_{j+1,x}$, $\tilde{K}=\Theta^{\prime}_{j+1,x}$ and $L_{K}=L_{j+1,x}$, it follows that $v_{t}^{\omega}\geq\beta/L_{j+1,x}$ for all $t\in\{1,\dots,T_{\theta_{0},B}^{\omega}\}$. From this result, we readily obtain that
$C^{\omega}_{\theta_{0},B,t}\leq D_{j}n_{B},\quad D_{j}=(\beta^{k_{j}}+1)(C_{\mathrm{eval}}+C_{\mathrm{grad}}),\quad k_{j}=\inf\{k\in\mathbb{N}:\,\beta^{k}<\beta/L_{j+1,x}\}.$